# Provably Extracting the Features from a General Superposition

### v1

@@ Abstract
It is widely believed that complex machine learning models generally encode features through linear representations, but these features exist in superposition, making them challenging to recover. We study the following fundamental setting for learning features in superposition from black-box query access: we are given query access to a function
$f(x)=\sum_{i=1}^{n}a_{i}\,\sigma_{i}(v_{i}^{\top}x),$
@@
In learning-theoretic terms, superposition refers to the overcomplete regime, when the number of features is larger than the underlying dimension (i.e. $n>d$), which has proven especially challenging for typical algorithmic approaches. Our main result is an efficient query algorithm that, from noisy oracle access to $f$, identifies all feature directions whose responses are non-degenerate and reconstructs the function $f$. Crucially, our algorithm works in a significantly more general setting than all related prior results — we allow for essentially arbitrary superpositions, only requiring that $v_{i},v_{j}$ are not nearly identical for $i\neq j$, and general response functions $\sigma_{i}$. At a high level, our algorithm introduces an approach for searching in Fourier space by iteratively refining the search space to locate the hidden directions $v_{i}$.
@@ 1 Introduction
While modern machine learning models are incredibly complex, a foundational viewpoint in our quest to understand them is the notion of linear representations [alain2016understanding, elhage2022toy, park2023linear, olah_linear_representation_2024, golowich2025sequenceslogitsreveallow]. The hypothesis is that salient features correspond to directions $v\in{\mathbb{R}}^{d}$ in some representation space, and the response to a feature depends only on the one–dimensional projection $v^{\top}x$ via some ridge function say $f(x)=\sigma(v^{\top}x)$.
Even the study of simple, single-feature functions, that depend only on the projection of the input onto a single direction, has led to a rich body of work in computational learning theory through generalized linear models [mccullagh2019generalized, kakade2011efficient, chen2020classification], single-index models [ichimura1993semiparametric, dudeja2018learning, bietti2022learning, gollakota2023agnostically, damian2023smoothing, damian2024computational, zarifis2024robustly], and also related problems such as non-Gaussian component analysis [blanchard2006search, diakonikolas2022non]. However, further adding to the challenge is the fact that for most models or functions that we want to study, the output depends on many features of the input. We formalize this in the most basic setting, when the output is a sum of the responses to individual features i.e. a sum of ridge functions
$f(x)\;=\;\sum_{i=1}^{n}a_{i}\,\sigma_{i}(v_{i}^{\top}x),$
@@
A fundamental concept in feature learning and interpretability is the notion ofsuperposition [elhage2022toy], meaning that the number of features is generally much larger than the dimension of the representation. In learning-theoretic terms, this corresponds to the overcomplete regime where the number of features $n$ exceeds the ambient dimension $d$. Yet, this overcomplete regime has proven particularly challenging from an algorithmic perspective, posing a barrier for common approaches such as moment and tensor methods [hillar2013most, anandkumar2013overcomplete, anandkumar2014tensor].
To complete the problem setup, we need to specify how we can access the function $f$. Typical learning setups assume $x$ is drawn from some distribution, such as a Gaussian, and then ask to recover $f$ given polynomially many samples $(x,f(x))$. However, it is difficult to posit tractable but realisic distributional assumptions — common assumptions like $x$ being Gaussian are unrealistic — and furthermore, even then there are still computational barriers in simple settings with just a single feature [goel2020statistical, song2021cryptographic, damian2024computational]. Instead, we study a query learning model where we assume black-box query access to some function $f_{\sim}$ with $\|f_{\sim}-f\|_{\infty}\leq\varepsilon$. While this is a stronger access model, it is natural from the lens of extracting features from a trained model. This query model also captures settings for model distillation and stealing, where a learner tries to recover something about a trained machine learning model through black-box API access [tramer2016stealing, finlayson2024logits, carlini2024stealing]. We now ask:
@@
Given queries to a function that is a sum of features, can we learn the underlying features? Are there barriers to learning if the features exhibit superposition?
This question is important both for understanding how to extract interpretable features (see e.g. [agarwal] which uses a special case of our model for interpretable deep learning) and from a fundamental computational learning theory perspective, where it builds on a vast existing body of work on learning single hidden layer neural networks. Within this literature there are two main points of comparison. First, there are numerous results in the standard “passive” setting, but the bottom line is that here, there is substantial evidence of computational hardness. There are statistical query lower bounds even for settings that are much simpler, special cases of our setting such as learning a single ridge function with general activation [goel2020statistical] and learning a linear combination of many ridge functions with fixed e.g. ReLU activation [diakonikolas2020algorithms, goel2020superpolynomial], suggesting that a general learning guarantee is out of reach for efficient algorithms. In query learning models, there has been work specifically in the case of ReLU neural networks i.e. $\sigma_{i}(z)=\text{ReLU}(z)$ for all $i\in[n]$, but these algorithms are extremely tailored to specifically the ReLU activation function, relying on trying to find the kink in the ReLU [rolnick2020reverse, chen2021efficientlylearninghiddenlayer, daniely2021exact].
@@
Our main result is an algorithm for learning a general sum of features whose query complexity and runtime is polynomial in all relevant parameters that, under mild and information-theoretically necessary assumptions — nontrivial separation between the feature directions $v_{i}$ and Lipschitzness of the functions $\sigma_{i}(\cdot)$ (see 1) — achieves the following guarantees:
@@
We emphasize that our result holds for overcomplete $v_{i}$, even when some of the $v_{i}$ are highly correlated and for general response functions $\sigma_{i}$. This circumvents the aforementioned computational hardness in the passive setting and significantly generalizes previous learning results which either require restrictive assumptions on the $\sigma_{i}$ (e.g. ReLU) [chen2021efficientlylearninghiddenlayer, daniely2021exact] or the $v_{i}$ (e.g. linear independence or near-orthogonality) [sedghi2016provable, oko2024learning].
@@
We also highlight the importance of the identifiability aspect of our result from the lens of feature learning and interpretability. As formulated in [elhage2022toy], one of the fundamental challenges in extracting features from modern models, and interpretability more broadly, is superposition — when the number of features is larger than the ambient dimension, it makes the features nonidentifiable. Indeed, this would be the case if the response functions were linear. However, from this viewpoint, our results actually give a new reason for optimism. The features are identifiable in a superposition as long as the response functions are nonlinear.
@@
Key to our result is a different type of algorithmic approach that departs from the typical method-of-moments recipe that is ubiquitous for learning latent variable models, especially over Euclidean spaces. The starting point is the basic observation that the Fourier transform $\widehat{f}$ of $f$ is “sparse”. Ignoring integrability issues for now, for a single ridge function $f(x)=\sigma(v^{\top}x)$, the Fourier transform is nonzero only on the line $\{tv\}_{t\in{\mathbb{R}}}$. Thus when $f$ is a sum of ridge functions, its Fourier transform is supported on only the lines through the origin in the directions $v_{i}$. Then, to recover the directions, we design an algorithm that can “locate the mass” in Fourier space. This paradigm has already been successful in boolean [goldreich1989hard, kushilevitz1991learning] and discrete [hassanieh2012nearly] settings. Our algorithm draws inspiration from this approach — the main subroutine (Algorithm 2) involves an iterative algorithm for searching in Fourier space to locate the hidden directions. However, compared to discrete spaces, the continuity and unboundedness of Euclidean space pose technical challenges — we emphasize the conceptual novelty here, that through carefully chosen geometric constructions and filter functions, we develop an algorithm that efficiently searches over a high-dimensional Euclidean Fourier domain. We provide a more detailed overview of our techniques in Section 3.
@@ 1.1 Related Work
There is a substantial body of work on GLM/SIM estimators, which recover a function that depends only on the projection of the input onto one direction, under various distributional and noise assumptions [mccullagh2019generalized, ichimura1993semiparametric, kakade2011efficient, dudeja2018learning, bietti2022learning, damian2023smoothing, gollakota2023agnostically, zarifis2024robustly, damian2024computational]. Multi-index models (MIMs) generalize SIMs to functions that depend only the projection of the input onto a constant dimensional subspace. Recently there has also been a growing body of work on learning MIMs [bietti2023learning, diakonikolas2024agnostically, mousavi2024learning]. We refer the reader to [bruna2025survey] for a more detailed overview of this literature. Note that our setting, involving a linear combination of ridge functions, is very different because the function actually depends on the full $d$-dimensional space.
@@
In our setting, when all of the $\sigma_{i}$ are some fixed activation function such as sigmoid or ReLU, then $f$ can be viewed as a two-layer neural network. There is a large body of work on learning two-layer neural networks from samples, which aims to characterize the boundaries of efficient algorithms and computational hardness e.g. [awasthi2021efficient, diakonikolas2020algorithms, ge2018learning, goel2020superpolynomial, chen24faster]. Beyond understanding the computational landscape, there is also a line of work towards understanding specifically the dynamics of gradient descent for training shallow networks e.g. [janzamin2016beatingperilsnonconvexityguaranteed, ge2017learningonehiddenlayerneuralnetworks, zhong2017recoveryguaranteesonehiddenlayerneural]. However in the passive setting, there are exponential in $\min(d,n)$ SQ lower bounds, suggesting that a general, efficient (i.e. $\mathrm{poly}(d,n)$) time algorithm is not possible.
There has also been work on learning neural networks with query access [rolnick2020reverse, chen2021efficientlylearninghiddenlayer, daniely2021exact]. However, as mentioned earlier, these results are extremely tailored to ReLU activations, whereas our result allows for arbitrary activations $\sigma_{i}$.
Linear combinations of more general activation functions are considered in [sedghi2016provable, oko2024learning]. These works also study a passive learning setting, and thus require structural assumptions on the $v_{i}$, namely that they are linearly independent in the former, and nearly orthogonal in the latter, whereas we do not need any such conditions.
@@
Our algorithm of searching in a high-dimensional Fourier space draws inspiration from the Goldreich–Levin heavy‑Fourier‑coefficient search algorithm [goldreich1989hard, kushilevitz1991learning] and also bears resemblance to algorithms used for computing sparse Fourier transforms [hassanieh2012nearly]. However, a key difference is that we give an efficient algorithm for searching over a high-dimensional Euclidean space. We remark that the connection between Fourier sparsity and expressing functions as a sum of ridge functions dates back to the classical works [barron, candes, donoho] but our focus here is on making this connection algorithmic.
@@
With the rise of APIs that provide users with black-box access to trained machine learning models, there has been increased interest in understanding what we can extract from such an API, both from a theoretical [mahajan2023learning, liu2025model, golowich2025sequenceslogitsreveallow, golowich2025provablylearningmodernlanguage] and practical perspective [tramer2016stealing, finlayson2024logits, carlini2024stealing]. Our result says that if a model is close to a sum of features in some representation space that we know, then we can recover the model from queries.
@@ 2.1 Ridge Functions (Features)
$f(x)=a_{1}\sigma_{1}(v_{1}^{\top}x)+\dots+a_{n}\sigma_{n}(v_{n}^{\top}x)\,.$
@@ 2.2 Learning Setup
$f(x)=a_{1}\sigma_{1}(v_{1}^{\top}x)+\dots+a_{n}\sigma_{n}(v_{n}^{\top}x)\,.$
@@
$\widetilde{f}(x)=a_{1}^{\prime}\sigma_{1}^{\prime}({v_{1}^{\prime}}^{\top}x)+\dots+a_{n}^{\prime}\sigma_{n}^{\prime}({v_{n}^{\prime}}^{\top}x)$
@@ 3 Technical Overview
Recall that the starting point is the intuition that if $f$ is a sum of $n$ features in directions $v_{1},\dots,v_{n}$, then the Fourier transform $\widehat{f}$ is nonzero only on the lines $\{tv_{i}\}_{t\in{\mathbb{R}}}$. First, to make this formal, we need to resolve the integrability issues. We define the Gaussian-reweighted function
@@
and we will choose $\ell$ sufficiently large. This ensures integrability and also that the Fourier transform remains concentrated around the lines $\{tv_{i}\}_{t\in{\mathbb{R}}}$. For a single ridge function $\rho(x)=\sigma(v^{\top}x)$, the Fourier transform can be written as (see 4.11)
@@
i.e., it concentrates on a tube of width $\approx 1/\ell$ around the line $\{tv:t\in{\mathbb{R}}\}$. A sum of ridge functions thus yields a sum of such tubes in Fourier space.

### v2

@@ Abstract
It is widely believed that complex machine learning models generally encode features through linear representations. This is the foundational hypothesis behind a vast body of work on interpretability. A key challenge toward extracting interpretable features, however, is that they exist in superposition. In this work, we study the question of extracting features in superposition from a learning theoretic perspective. We start with the following fundamental setting: we are given query access to a function
$f(x)=\sum_{i=1}^{n}\sigma_{i}(v_{i}^{\top}x),$
@@
In learning-theoretic terms, superposition refers to the overcomplete regime, when the number of features is larger than the underlying dimension (i.e. $n>d$), which has proven especially challenging for typical algorithmic approaches. Our main result is an efficient query algorithm that, from noisy oracle access to $f$, identifies all feature directions whose responses are non-degenerate and reconstructs the function $f$. Crucially, our algorithm works in a significantly more general setting than all related prior results. We allow for essentially arbitrary superpositions, only requiring that $v_{i},v_{j}$ are not nearly identical for $i\neq j$, and allowing for general response functions $\sigma_{i}$. At a high level, our algorithm introduces an approach for searching in Fourier space by iteratively refining the search space to locate the hidden directions $v_{i}$.
@@ 1 Introduction
While modern machine learning models are incredibly complex, a foundational viewpoint in our quest to understand them is the notion of linear representations [2, 28, 51, 50, 37]. The hypothesis is that salient features correspond to directions $v\in{\mathbb{R}}^{d}$ in some representation space, and the response to a feature depends only on the one–dimensional projection $v^{\top}x$ via some ridge function say $f(x)=\sigma(v^{\top}x)$. This hypothesis has served as the basis for a vast body of work on extracting features, activation steering, and interpretability more broadly [43, 56, 50].
From a theoretical perspective, even the study of simple, single-feature functions, that depend only on the projection of the input onto a single direction, has led to a rich body of work in computational learning theory through generalized linear models [47, 42, 15], single-index models [40, 25, 8, 35, 18, 19, 57], and also related problems such as non-Gaussian component analysis [9, 23]. However, further adding to the challenge is the fact that for most models or functions that we want to study, the output depends on many features of the input. We formalize this in the most basic setting, when the output is a sum of the responses to individual features i.e. a sum of ridge functions
$f(x)\;=\;\sum_{i=1}^{n}\sigma_{i}(v_{i}^{\top}x),$
@@
A key concept in feature learning and interpretability is the notion of superposition, formalized in [28]. Superposition refers to the fact that the number of features is generally much larger than the dimension of the representation, and has been highlighted as an important obstacle and potentially fundamental barrier to extracting interpretable features [28, 50, 17]. As stated in [27], “superposition means there is no basis in which activations are interpretable”. This means that we cannot obtain an ideal decomposition into features that are disentangled and correspond to human-interpretable concepts.
In practice, sparse autoencoders and related dictionary-learning methods have been used to recover interpretable features from representations in superposition in some cases [10, 17]. However, these empirical methods are optimization heuristics and can exhibit instability, incompleteness, and overfitting [50], while theoretical guarantees apply only under strong structural assumptions on the underlying features. This motivates us to ask the following question from a learning-theoretic perspective:
Guiding Question.
Can we study superposition and the obstacles it poses from the perspective of computational learning theory? Can we formalize the challenges that arise in simple settings and develop algorithms to address them?
In learning-theoretic terms, superposition corresponds to the overcomplete regime where the number of features $n$ exceeds the ambient dimension $d$. Yet, this overcomplete regime has also proven particularly challenging from an algorithmic perspective, posing a barrier for common approaches such as moment and tensor methods [39, 3, 4].
To formalize the problem setup, it remains to specify how we can access the function $f$. From the perspective of extracting features from a trained model, the natural question is to recover the features given the ability to query the function. In fact, this is the typical setting for interpretability analyses, where one takes a trained model and then analyzes its activations and tries to understand the learned representations by plugging in carefully chosen inputs. Thus, we study the following query learning setting: we assume black-box query access to some function $f_{\sim}$ with $\|f_{\sim}-f\|_{\infty}\leq\varepsilon$. We now ask:
@@
Given queries to a function that is (close to) a sum of features, can we learn the underlying features? Are there barriers to learning if the features exhibit superposition?
Beyond being natural from the perspective of feature learning and interpretability, this query learning model also allows us to sidestep the difficulty of positing realistic distributional assumptions. Typical learning setups often assume $x$ is drawn from some distribution, such as a Gaussian, and then ask to recover $f$ given polynomially many samples $(x,f(x))$. However, it is difficult to posit tractable but realistic distributional assumptions — common assumptions like $x$ being Gaussian are unrealistic — and furthermore, even then there are still computational barriers in simple settings with just a single feature [33, 54, 19]. The query model also captures natural settings for model distillation and stealing, where a learner tries to recover something about a trained machine learning model through black-box API access [55, 29, 13].
In light of the discussion above, the main question we pose is important both for understanding how to extract interpretable features  [^2] Also see e.g. [1] which proposes Neural Additive Models (NAMs), a special case of our model, as an architecture for interpretable deep learning and from a fundamental computational learning theory perspective, where it builds on a vast existing body of work on learning single hidden layer neural networks. Within this literature there are two main points of comparison. First, there are numerous results in the standard “passive” setting, but the bottom line is that here, there is substantial evidence of computational hardness. There are statistical query lower bounds even for settings that are much simpler, special cases of our setting such as learning a single ridge function with general activation [33] and learning a linear combination of many ridge functions with fixed e.g. ReLU activation [22, 32], suggesting that a general learning guarantee is out of reach for efficient algorithms. Second, in query learning models, there has been work specifically in the case of ReLU neural networks i.e. $\sigma_{i}(z)=\text{ReLU}(z)$ for all $i\in[n]$, but these algorithms are extremely tailored to specifically the ReLU activation function, relying on trying to find the kink in the ReLU [52, 14, 20]. From the lens of feature learning, we expect that response functions can be far more general than just ReLU, and thus previous work falls significantly short of a clean resolution.
@@
Our main result is an algorithm for learning a general sum of features under mild and information-theoretically necessary assumptions — nontrivial separation between the feature directions $v_{i}$ and Lipschitzness of the functions $\sigma_{i}(\cdot)$ (see 1). Our algorithm has query complexity and runtime polynomial in all relevant parameters and achieves the following guarantees:
@@
We emphasize that our result holds for overcomplete $v_{i}$, even when some of the $v_{i}$ are highly correlated and for general response functions $\sigma_{i}$. This circumvents the aforementioned computational hardness in the passive setting and significantly generalizes previous learning results which either require restrictive assumptions on the $\sigma_{i}$ (e.g. ReLU) [14, 20] or the $v_{i}$ (e.g. linear independence or near-orthogonality) [53, 49].
@@
We also highlight the importance of the identifiability aspect of our result from the lens of feature learning and interpretability. As formulated in [28], one of the fundamental challenges in extracting features from modern models, and interpretability more broadly, is superposition — when the number of features is larger than the ambient dimension, it makes the features nonidentifiable. Indeed, this would be the case if the response functions were linear. However, from this viewpoint, our results actually give a new reason for optimism. A key takeaway from our results is:
Takeaway.
For a function that is a sum of many features, the individual features are identifiable, even in a superposition, as long as the response functions are nonlinear.
@@
Key to our result is a different type of algorithmic approach that departs from the typical method-of-moments recipe that is ubiquitous for learning latent variable models, especially over Euclidean spaces. The starting point is the basic observation that the Fourier transform $\widehat{f}$ of $f$ is “sparse”. Ignoring integrability issues for now, for a single ridge function $f(x)=\sigma(v^{\top}x)$, the Fourier transform is nonzero only on the line $\{tv\}_{t\in{\mathbb{R}}}$. Thus when $f$ is a sum of ridge functions, its Fourier transform is supported on only the lines through the origin in the directions $v_{i}$. This observation was made in [26] and has been used for proving negative results and depth separations about the expressibility of shallow neural networks. However, here we will use this structural property algorithmically for a positive result.
At a high level, to recover the directions, we design an algorithm that can “locate the mass” in Fourier space. This paradigm has already been successful in boolean [34, 44] and discrete [38] settings. Our algorithm draws inspiration from this approach — the main subroutine (Algorithm 2) involves an iterative algorithm for searching in Fourier space to locate the hidden directions. However, compared to discrete spaces, the continuity and unboundedness of Euclidean space pose technical challenges — we emphasize the conceptual novelty here, that through carefully chosen geometric constructions and filter functions, we develop an algorithm that efficiently searches over a high-dimensional Euclidean Fourier domain. We provide a more detailed overview of our techniques in Section 3.
@@ 1.1 Related Work
There is a substantial body of work on GLM/SIM estimators, which recover a function that depends only on the projection of the input onto one direction, under various distributional and noise assumptions [47, 40, 42, 25, 8, 18, 35, 57, 19]. Multi-index models (MIMs) generalize SIMs to functions that depend only on the projection of the input onto a constant dimensional subspace. Recently there has also been a growing body of work on learning MIMs [7, 21, 48]. We refer the reader to [11] for a more detailed overview of this literature. Note that our setting, involving a linear combination of ridge functions, is very different because the function actually depends on the full $d$-dimensional space.
@@
In our setting, when all of the $\sigma_{i}$ are some fixed activation function such as sigmoid or ReLU, then $f$ can be viewed as a two-layer neural network. There is a large body of work on learning two-layer neural networks from samples, which aims to characterize the boundaries of efficient algorithms and computational hardness e.g. [5, 22, 30, 32, 16]. Beyond understanding the computational landscape, there is also a line of work towards understanding specifically the dynamics of gradient descent for training shallow networks e.g. [41, 31, 58]. However in the passive setting, there are exponential in $\min(d,n)$ SQ lower bounds, suggesting that a general, efficient (i.e. $\mathrm{poly}(d,n)$) time algorithm is not possible.
There has also been work on learning neural networks with query access [52, 14, 20]. However, as mentioned earlier, these results are extremely tailored to ReLU activations, whereas our result allows for arbitrary activations $\sigma_{i}$.
Linear combinations of more general activation functions are considered in [53, 49]. These works also study a passive learning setting, and thus require structural assumptions on the $v_{i}$, namely that they are linearly independent in the former, and nearly orthogonal in the latter, whereas we do not need any such conditions.
@@
Our algorithm of searching in a high-dimensional Fourier space draws inspiration from the Goldreich–Levin heavy‑Fourier‑coefficient search algorithm [34, 44] and also bears resemblance to algorithms used for computing sparse Fourier transforms [38]. However, a key difference is that we give an efficient algorithm for searching over a high-dimensional Euclidean space. We remark that the connection between Fourier sparsity and expressing functions as a sum of ridge functions dates back to the classical works [6, 12, 24] but our focus here is on making this connection algorithmic.
@@
With the rise of APIs that provide users with black-box access to trained machine learning models, there has been increased interest in understanding what we can extract from such an API, both from a theoretical [46, 45, 37, 36] and practical perspective [55, 29, 13]. Our result says that if a model is close to a sum of features in some representation space that we know, then we can recover the model from queries.
@@ 2.1 Ridge Functions (Features)
$f(x)=\sigma_{1}(v_{1}^{\top}x)+\dots+\sigma_{n}(v_{n}^{\top}x)\,.$
Note that the terms sum and linear combination can be used interchangeably since we can always absorb the coefficients of a linear combination into the functions $\sigma_{i}$ themselves.
@@ 2.2 Learning Setup
$f(x)=\sigma_{1}(v_{1}^{\top}x)+\dots+\sigma_{n}(v_{n}^{\top}x)\,.$
@@
$\widetilde{f}(x)=\sigma_{1}^{\prime}({v_{1}^{\prime}}^{\top}x)+\dots+\sigma_{n}^{\prime}({v_{n}^{\prime}}^{\top}x)$
@@
2.3 Main Results
The final theorem and corresponding identifiability result are stated below. After completing the proof of Theorem 2.5, we complete the proof of the stronger version through a direct reduction in Section 8.2.
Theorem 2.7.
$\widetilde{f}(x)=\sigma_{1}^{\prime}({v_{1}^{\prime}}^{\top}x)+\dots+\sigma_{n}^{\prime}({v_{n}^{\prime}}^{\top}x)$
Corollary 2.8.
@@ 3 Technical Overview
Recall that the starting point is the intuition that if $f$ is a sum of $n$ features in directions $v_{1},\dots,v_{n}$, then the Fourier transform $\widehat{f}$ is nonzero only on the lines $\{tv_{i}\}_{t\in{\mathbb{R}}}$. First, to make this formal, we need to resolve the integrability issues. Recall the Fourier transform is defined as
$\widehat{f}(y)=\frac{1}{(2\pi)^{d/2}}\int_{{\mathbb{R}}^{d}}e^{-\mathrm{i}y^{\top}x}f(x)dx$
but for the above to be well-defined, we require $f(x)$ to be integrable over ${\mathbb{R}}^{d}$. To address this, we will multiply by a Gaussian reweighting, which will ensure integrability. We define the Gaussian-reweighted function
@@
and we will choose $\ell$ sufficiently large. With this definition, it is clear that the Fourier transform $\widehat{f^{(\ell)}}$ is well-defined. Furthermore, it remains concentrated around the lines $\{tv_{i}\}_{t\in{\mathbb{R}}}$. To see why, for a single ridge function $\rho(x)=\sigma(v^{\top}x)$, the Fourier transform can be written as (see 4.11)
@@
and the exponential term implies that it concentrates on a tube of width $\approx 1/\ell$ around the line $\{tv:t\in{\mathbb{R}}\}$. A sum of ridge functions thus yields a sum of such tubes in Fourier space.